# Uncovering Reliable Indicators: Improving IoC Extraction from Threat Reports

### v1

@@ 5.2.2 Timing Evaluation
From the results of those measurements, we see a 43% drop in the average and median time the analysts spent on a report. These findings validate a key goal of our HITL approach: significantly reducing analyst workload without compromising accuracy. The 43% reduction in annotation time, when paired with consistently high precision and recall, demonstrates that LLM-augmented interfaces can streamline expert workflows by offloading routine labeling while still allowing humans to oversee critical edge cases.
@@ Appendix A LANCE Prompts
This section presents the prompt engineering phase for LANCE as well as the evaluation of the final prompts. For both the prompt engineering phase and the evaluation of the prompts, we used ChatGPT 4o [52] as the underlying LLM.
@@ A.1 Prompt Engineering and Input Structuring
TABLE IV: LANCE prompt structure and example. The example prompt in the corresponding column is the final prompt used for LANCE for the labeling of URLs. The same techniques and structures were used as needed for the rest of the IoC Types.
To evaluate prompt performance, we manually extracted and labeled all indicators in 10 reports. These reports were selected from AlienVault with the turgent of ensuring diversity in structure, size, and IoC density. The indicator extraction and labeling process was iterative, as some indicators were missed or misclassified in the initial extraction iteration.
@@
Due to the nature of the different types of indicators, as well as the reasons for appearing in a report, some types were more difficult to prompt than others. Specifically, domain labeling was the most difficult task to prompt-engineer for. In our experiments, we came across several domain-like strings that were not actual domains. This was to be expected if taking into consideration the analysis in Section 2.3. In addition, the appearance of nonIoC domains in IoC URLs was very common, making it necessary for the prompt to cover these instances of domains.
@@
The techniques analyzed in Section 3.2.2 as well as the corresponding parts from the final prompt used for URLs, can be seen in Table IV. By constructing our prompts this way, we address most of the pitfalls the LLM can fall into and steer it away from common inaccuracies.
@@ Appendix B LANCE Generalizability
TABLE V: Comparison of the LANCE implementation using GPT, Llama, Nvidia Nemotron, Gemini, and Gemma on IoC Classification.
In Table V we see a detailed evaluation of LANCE using 5 state of the art LLMs for the LLM component: ChatGPT 4o [52], LLama 3.3 70b [56], Nvidia Llama 3.1 Nemotron 70b [59], Gemini 2.0 Flash [58], and Gemma 3 27b [57]. The table shows the precision, recall, and F1 score the pipeline achieved with each model per indicator type and in total. The total F1 score for each LLM is also shown in Figure 9.
Appendix C Ambiguous IoCs’ Labels
In Figure 11 we show a segment of a report which is part of PRISM dataset. Several URLs of legitimate websites were compromised and used as command and control (C2) points for the campaign. According to our senior analyst, the domains (e.g., “abert-online.de”) are not IoCs (as they are a legitimate website), whereas the complete URLs are. LANCE labeled these domains as “IoC” and justified the label(see Table VI) by referencing their “involvement in the malicious activities.” Notably, LANCE acknowledged the domains’ legitimate nature but still marked them as IoCs due to their role in the attack. The justifications for the remaining domains in Figure 11 were similar, and all were consistently labeled as IoCs. In this case, both interpretations – labeling as IoC or non-IoC – can be considered correct, depending on the downstream use of the extracted indicators.
Figure 11: Segment of a report included in PRISM
LANCE justification
The domain abert-online.de is classified as an IoC because it
is identified as a legitimate website that has been compromised
and used as a command and control (C2) server for JackalControl,
indicating its involvement in malicious activities.
TABLE VI: The justification LANCE gave for the “IoC” label it assigned to domain “abert-online.de” which appear in the report segment shown in Figure 11.
@@ Appendix D Ground truth creation methods evaluation details
In Table VII we see the performance of the five ground truth creation methods discussed in Section 5.1.1 on coverage as well as Precision, Recall, and F1 score on each type of IoCs (IPs, hashes, URLs, and domains).
@@
TABLE VII: Performance comparison of four prominent ground truth creation methods from Table I, across four IoC types (IP, hash, URL, domain). The RegEx + Whitelists methodology [1] exhibits near-perfect recall but low precision due to excessive false positives. VirusTotal-based methods vary depending on the chosen threshold, but they have a high number of unlabeled indicators. The same can be said for AlienVault, which has the highest number of unlabeled indicators. LANCE achieves the highest F1 score across all indicator types while providing excellent coverage.
@@ Appendix E User Interface
Figure 13: This is a view of the user interface, the analysts used for the BAP of the ground truth creation. The extracted Indicators can be seen marked in a blue rectangle, which changes color according to the label assigned by the analyst.

### v2

@@ 5.2.2 Timing Evaluation
From the results of those measurements, we see a $43\%$ drop in the average and median time the analysts spent on a report. These findings validate a key goal of our HITL approach: significantly reducing analyst workload without compromising accuracy. The $43\%$ reduction in annotation time, when paired with consistently high precision and recall, demonstrates that LLM-augmented interfaces can streamline expert workflows by offloading routine labeling while still allowing humans to oversee critical edge cases.
5.3 LANCE Misclassifications Analysis
In this section, we perform an analysis of common misclassifications of LANCE by discussing common cases of false positives and false negatives. These cases occurred when using ChatGPT-4o [55] as the backbone LLM. In Figure 10, we show a segment of a report which is part of PRISM dataset. Several domains and URLs of legitimate websites were compromised and used as command and control (C2) points for the campaign. According to our senior analyst, the domains (e.g., abert-online.de) are not IoCs (as they are a legitimate website), whereas the complete URLs are. LANCE labeled these domains as “IoC” and justified the label (see Table IV) by referencing their “involvement in the malicious activities.” Notably, LANCE acknowledged the domains’ legitimate nature but still marked them as IoCs due to their role in the attack. The justifications for the remaining domains in Figure 10 were similar, and all were consistently labeled as IoCs. In this case, both interpretations (IoC or nonIoC) can be considered correct, depending on the downstream use of the extracted indicators. These mislabeled cases highlight a key challenge: the definition of what constitutes an IoC can vary depending on the downstream task. For example, a compromised legitimate domain might be labeled as an IoC for detection or blocking purposes, but as nonIoC in datasets intended for long-term attribution or infrastructure tracking. To mitigate this ambiguity, systems should explicitly define IoC vs. nonIoC labels in relation to their intended use case and surface this definition to analysts during labeling. Such task-aware definitions, combined with LANCE’s explanatory justifications, can reduce false negatives by helping analysts resolve edge cases consistently and in alignment with the downstream application.
Figure 10: Segment of a report included in PRISM
LANCE justification
The domain abert-online.de is classified as an IoC because it
is identified as a legitimate website that has been compromised
and used as a command and control (C2) server for JackalControl,
indicating its involvement in malicious activities.
TABLE IV: The justification LANCE gave for the “IoC” label it assigned to domain abert-online.de which appear in the report segment shown in Figure 10.
Beyond the ambiguous cases of compromised but legitimate domains, we observed two recurring categories of false negative (FN) domain labels. First, Firebase subdomains (e.g., domain.firebase.com) were often mislabeled as nonIoCs because the LLM defaulted to seeing Firebase as benign despite malicious context. These cases illustrate the tension between platform ubiquity and malicious abuse, and the need for downstream-task-aware definitions to guide consistent labeling.
Second, a systematic error emerged with domain variants containing the www. prefix. Indicators presented as www.domain.com were labeled nonIoC, while the same domain without the prefix was correctly labeled as IoC. Interestingly, in these cases, the LLM justifications often explicitly stated that the www. variant was malicious, even when the final label output was nonIoC. This disconnect highlights the value of explainable justifications that allowed analysts to override errors during HITL validation.
@@ Appendix A LANCE Prompts
This section presents the prompt engineering phase for LANCE as well as the evaluation of the final prompts. For both the prompt engineering phase and the evaluation of the prompts, we used ChatGPT-4o [55] as the underlying LLM.
@@ A.1 Prompt Engineering and Input Structuring
TABLE V: LANCE prompt structure and example. The example prompt in the corresponding column is the final prompt used for LANCE for the labeling of URLs. The same techniques and structures were used as needed for the rest of the IoC Types.
To evaluate prompt performance, we manually extracted and labeled all indicators in 10 reports. These reports were selected from AlienVault with the intent of ensuring diversity in structure, size, and IoC density. The indicator extraction and labeling process was iterative, as some indicators were missed or misclassified in the initial extraction iteration.
@@
Due to the nature of the different types of indicators, as well as the reasons for appearing in a report, some types were more difficult to prompt than others. Specifically, domain labeling was the most difficult task to prompt-engineer. In our experiments, we came across several domain-like strings that were not actual domains. This was to be expected if taking into consideration the analysis in Section 2.3. In addition, the appearance of nonIoC domains in IoC URLs was very common, making it necessary for the prompt to cover these instances of domains.
@@
The techniques analyzed in Section 3.2.2, as well as the corresponding parts from the final prompt used for URLs, can be seen in Table V. By constructing our prompts this way, we address most of the pitfalls the LLM can fall into and steer it away from common inaccuracies.
@@ A.2 Prompt Evaluation
TABLE VI: Comparison of the LANCE implementation using GPT, Llama, Nvidia Nemotron, Gemini, and Gemma on IoC Classification.
Appendix B Evaluation of IoC extraction components of CTI extraction tools
Table VII presents the precision, recall, and F1 scores of the IoC extraction components of representative prior systems that focus on CTI extraction on PRISM. Extractor [38], LADDER [39], and CTINexus [40] were originally designed for broader CTI extraction tasks such as attack pattern and knowledge graph construction. Their indicator extraction components rely primarily on regular expressions and are therefore not optimized for precise IoC labeling. This limitation is evident in their relatively low F1 scores on PRISM, with CTINexus in particular achieving very poor recall. IoCSearcher [1], a state-of-the-art RegEx tool, attains perfect recall but at the cost of precision. In contrast, LANCE balances both dimensions, achieving an F1 score of 0.98. These results underscore that while prior CTI systems contribute valuable higher-level intelligence, reliable IoC extraction and ground truth creation require special attention. Moreover, integrating LANCE into such systems could substantially improve their IoC extraction performance, providing them with a more accurate foundation on which to build higher-level CTI analysis.
Tool Precision Recall F1
IoCSearcher [1] 0.44 1.00 0.61
Extractor [38] 0.38 0.32 0.35
LADDER [39] 0.70 0.37 0.48
CTINexus REGEX only [40] 0.30 0.33 0.31
CTINexus [40] 0.52 0.02 0.03
LANCE 0.98 0.98 0.98
TABLE VII: Precision, Recall, and F1 score for IoC extraction components of CTI extraction tools on PRISM.
Appendix C LANCE Generalizability
In Table VI we see a detailed evaluation of LANCE using 5 state of the art LLMs for the LLM component: ChatGPT 4o [55], LLama 3.3 70b [59], Nvidia Llama 3.1 Nemotron 70b [62], Gemini 2.0 Flash [61], and Gemma 3 27b [60]. The table shows the precision, recall, and F1 score the pipeline achieved with each model per indicator type and in total. The total F1 score for each LLM is also shown in Figure 9.
@@ Appendix D Ground truth creation methods evaluation details
In Table VIII we see the performance of the five ground truth creation methods discussed in Section 5.1.1 on coverage as well as Precision, Recall, and F1 score on each type of IoCs (IPs, hashes, URLs, and domains).
@@
TABLE VIII: Performance comparison of four prominent ground truth creation methods from Table I, across four IoC types (IP, hash, URL, domain). The RegEx + Whitelists methodology [1] exhibits near-perfect recall but low precision due to excessive false positives. VirusTotal-based methods vary depending on the chosen threshold, but they have a high number of unlabeled indicators. The same can be said for AlienVault, which has the highest number of unlabeled indicators. LANCE achieves the highest F1 score across all indicator types while providing excellent coverage.
@@ Appendix E User Interface
Figure 13: This is a view of the user interface that the analysts used for the BAP of the ground truth creation. The extracted Indicators can be seen marked in a blue rectangle, which changes color according to the label assigned by the analyst.